OpenStudy (anonymous):

If six inscribed, wooden disks are cut from a 36 centimeter by 24 centimeter rectangular block of wood, how many square centimeters of wood will not be used?

OpenStudy (anonymous):

ill assume diameter is max

OpenStudy (anonymous):

i got 185.415

OpenStudy (dumbcow):

so 6d = 36 -> d = 6 then radius is 3 area = pi*3^2 multiply by 6 total area = pi*6*9 subtract that from rectangle area (36*24)

OpenStudy (lalaly):

i agree wit dumbcow

OpenStudy (lalaly):

and hahd xD

OpenStudy (anonymous):

24/2 and 36/3 gives max diameter

OpenStudy (anonymous):

yes so do i....both of you are right.....you guys are awsome....more still to come ;)

OpenStudy (dumbcow):

thats a lot of extra space, i dunno you could probably get bigger circles hahd is prob right

OpenStudy (anonymous):

haha it is right and you and lalay got the answer on the dot!!!